A stock will pay dividends of $1.20 four years from today, $1.50 five years from today, and $1.90 six years from today. There will be no dividend payments prior to year 4. After year 6, the growth rate in dividends per year is expected to be 6% forever. The required return on the stock is 11%. P4, the price of the stock 4 years from now, should be $_______.


* DO NOT ROUND INTERMEDIATE VALUES, NO CREDIT WILL BE GIVEN
* FINAL ANSWER IN DOLLARS, ROUNDED TO TWO DECIMAL PLACES
Answer #1

D5 = 1.5

D6 = 1.90

P6 = D7/(k-g) = 1.9*106%/(11%-6%)

= 40.28

At Year 4

P4 = D5/(1+k)+ (D6+P6) /(1+k)^2
= 1.5/ 1.11^1 + (1.9+40.28) /1.11^2

= $ 35.59

Hence price of the stock after 4 years = $ 35.59